A group of students held a car wash to raise money. They charged $9 for each car and $12 for each van or SUV. Two teams worked continuously, one washing cars and the other washing vans and SUVs. The team washing cars required 25% less time per vehicle than the other team. Each team washed vehicles for 4 hours. The two teams together raised a total of $720. How long, on average, did each team take to wash a vehicle?
Answer by richwmiller(17219)

9c+12v=720,
a=.75b,
ac=4,
bv=4
a=0.1 (6 minutes), b=0.133333 (8 minutes), c=40., v=30.
.1 (6 minutes) hour to wash a car and .1333 hours ( 8 minutes) to wash a van
30 cars and 40 vans
40*6=240 minutes
30*8=240
360+360=720